Calculate the density of the moon by assuming it to be a sphere of diameter 3475 km and having a mass of 7.35 × 1022 kg. Express your answer in g/cm3

1
Expert's answer
2022-06-02T10:26:06-0400
"\rho=\frac{m}{V}\\
=\frac{7.35*10^{22}\:\rm kg}{\frac{4}{3}*3.14*(3475*10^3/2)^3}\\
=3350\:\rm kg/m^3=3.35\: g/cm^3"